Analyzing Manufacturing Cost Accounts

Clapton Company manufactures custom guitars in a wide variety of styles. The following incomplete ledger accounts refer to transactions that are summarized for May:

Materials
May 1 Balance 37,500 May 31 Requisitions (a)
31 Purchases 150,600


Work in Process
May 1 Balance (b) 31 Completed jobs (f)
31 Materials (c)
31 Direct labor (d)
31 Factory overhead applied (e)


Finished Goods
May 1 Balance 0 May 31 Cost of goods sold (g)
31 Completed jobs (f)


Wages Payable
May 31 Wages incurred 121,300


Factory Overhead
May 1 Balance 21,700 May 31 Factory overhead applied (e)
31 Indirect labor (h)
31 Indirect materials 20,000
31 Other overhead 62,500

In addition, the following information is available:

  1. Materials and direct labor were applied to the following jobs in May:
    Job No. Style Quantity Direct Materials Direct Labor
    101 AF1 170 $21,030 $13,000
    102 AF3 430 44,180 27,000
    103 AF2 210 17,720 8,000
    104 VY1 300 45,900 30,000
    105 VY2 160 21,500 15,000
    106 AF4 150 9,850 4,000
    Total 1,420 $160,180 $97,000
  2. Factory overhead is applied to each job at a rate of 130% of direct labor cost.
  3. The May 1 Work in Process balance consisted of two jobs, as follows:
    Job No. Style Work in Process,
    May 1
    101 AF1 $5,000
    102 AF3 16,700
    Total $21,700
  4. Customer jobs completed and units sold in May were as follows:
    Job No. Style Completed in
    May
    Units Sold
    in May
    101 AF1 X 136
    102 AF3 X 344
    103 AF2 0
    104 VY1 X 252
    105 VY2 X 133
    106 AF4 0

Required:

1. Determine the missing amounts associated with each letter and complete the following table. If required, round amounts to the nearest dollar. If an answer is zero, enter in "0". Enter all amounts as positive numbers.

Job No. Quantity May 1
Work in
Process
Direct
Materials
Direct
Labor
Factory
Overhead
Total Cost Unit Cost Units Sold Cost of Goods Sold
No. 101 x $ 5,000 $ 21,030 $ 13,000 $x $x $x x $x
No. 102 x 16,700 44,180 27,000 x x x x x
No. 103 x 17,720 8,000 x x x x
No. 104 x 45,900 30,000 x x x x x
No. 105 x 21,500 15,000 x x x x x
No. 106 x 9,850 4,000 x x x x
Total x $21,700 $160,180 $97,000 $x $x $x

a. Materials Requisitions $__x_______

b. Work in Process Beginning Balance $_x________

c. Direct Materials $___x________

d. Direct Labor $____x______

e. Factory overhead applied $_____x_______

f. Completed jobs $___x________

g. Cost of goods sold $____x_______

h. Indirect labor $____x_________

2. Determine the May 31 balances for each of the inventory accounts and factory overhead. Use the minus sign to indicate any credit balances.

Materials $x
Work in Process $x
Finished Goods $x
Factory Overhead $x

Wherever there is an "x" theres an amount box for me to fill. Thanks!
Answer #1

Job No

Quantity

(A)

May 1 Work in process

Direct materials

Direct labor

Factory overhead

130% of direct labor cost

Total cost

(B)

Unit cost

(B)/(A)

Unit sold

Cost of goods sold

101

170

$5,000

$21,030

$13,000

$16,900

$55,930

$329

136

$44,744

102

430

16,700

44,180

27,000

35,100

122,980

286

344

98,384

103

210

17,720

8,000

10,400

36,120

172

0

0

104

300

45,900

30,000

39,000

114,900

383

252

96,516

105

160

21,500

15,000

19,500

56,000

350

133

46,550

106

150

9,850

4,000

5,200

19,050

127

0

0

Total

1,420

$21,700

$160,180

$97,000

$126,100

$404,980

$286,194

a. Material requisition = Total Direct material (as calculated above)+ Indirect material

= $160,180 + 20,000 = $180,180

b. Work in process beginning balance = $72,400 (as calculated above)

c. Direct material = $160,180

d. Direct labor = $97,000

e. Factory overhead applied = $126,100

f. Completed jobs = Job no. 101 + Job no. 102 + Job no. 104 + Job no. 105

= $55,930 + 122,980 + 114,900 + 56,000 = $349,810

g. Cost of goods sold = $286,194

h. Indirect labor = Wages incurred - Direct labor

= $121,300 - 97,000 = $24,300

2.

Computation of May 31 balances for each of the following:

Material = Beginning balance + purchases - requisition

= $37,500 + 150,600 - 180,180 = $7,920

Work in process = Job 103 + Job 106

= $36,120 + 19,050 = $55,170

Finished goods = Completed jobs - Cost of goods sold

= $349,810 - 286,194 = $63,616

Factory overhead =( Beginning balance + Indirect labor + indirect materials + other overhead ) - Factory overhead applied

=( $21,700 + 24,300 + 20,000 + 62,500) - 126,100

= $128,500 - 126,100

= $2,400

kindly give a ?. It helps me. Thanks!!